A bag contains 5 red balls and 4 blue balls. If 2 balls are selected at random, find the probability of selecting 2 red balls
Answer by greenestamps(13327)



This is a problem that is easily solved using two basic methods.

(1) Select the balls one at a time

P(1st ball is red) = 5/9 (there are 9 balls of which 5 are red)
P(2nd ball is red) = 4/8 (there are 8 balls left, of which 4 are red)
P(both balls are red) = (5/9)(4/8) = (5/9)(1/2) = 5/18

(2) Use "n choose r"

number of ways of choosing any 2 of the 9 balls: "9 choose 2" = (9*8)/(2*1) = 72/2 = 36
number of says of choosing 2 of the 5 red balls: "5 choose 2" = (5*4)/(2*1) = 20/2 = 10

P(choosing 2 of the red balls) = 10/36 = 5/18